# On Scaling LT-Coded Blockchains in Heterogeneous Networks and their Vulnerabilities to DoS Threats

### v2

@@ 4.1.2 Attack Strategy against OFG Decoder
[…]
OFG_Optimal_Attack A reasonable attack strategy against the OFG decoder involves the following steps:
@@
One possible method to carry out the aforementioned attack strategy is to undertake an exhaustive search to find the combination of droplet nodes that produces a matrix of minimum rank. However, this is a computationally expensive approach as it involves the evaluation of ranks of ($\sigma_{0}S\choose\sigma S$ matrices, each with dimensions $(\sigma_{0}-\sigma)S\times k$. To circumvent this problem, we present a computationally feasible algorithm, presented in Algorithm 1, as an alternative approach to exhaustive search. We emphasize that the algorithm may not identify the exact $(\sigma_{0}-\sigma)S$ droplet nodes of minimum rank, however, it does provide a solution with low rank.
@@
10:              Swap($\mathbf{R}[j][\cdot]$,$\mathbf{R}[\sigma_{0}S-i+2][\cdot]$);
@@ 4.1.3 Attack Strategy against BRH Decoder
[…]
BFG_Optimal_Attack A reasonable attack strategy against the BRH decoder is same as that against the OFG decoder.
@@ 4.1.4 Attack Strategy against CRH Decoder
[…]
@@
[…]
CRH_Optimal_Attack For the CRH decoder with $\eta_{c}>0$ , we propose to follow the same attack strategy as that against the BP decoder. However, for the CRH decoder with $\eta_{c}=0$ , we propose to follow the same attack strategy as that against the OFG decoder.
@@ 4.2 Experimental Results for Attack Strategies
Figure 10: Effectiveness of the proposed degree based attack strategy against a bucket node with BP decoder. The parameters for simulations are $k$=20, $S$=60, $\sigma\in\{0.1,0.2,0.3,0.4\}$ and $\xi\in\{1.3,1.6,1.9,2.2,2.5\}$.
[…]
Figure 11: Comparison of degree based and score based attack strategies against a bucket node with BP decoder. The parameters for simulations are $k$=20, $S$=60, $\sigma\in\{0.05,0.1,0.15,0.2\}$ and $\sigma_{0}\in\{0.6,0.7,0.8,0.9\}$.
Figure 12: Effectiveness of the proposed attack strategy against a bucket node with OFG decoder. The parameters for the simulations are $k$=20, $S$=60, $\sigma\in\{0.4,0.45,0.5,0.55\}$ and $\xi\in\{1.2,1.4,1.6,1.8\}$.
[…]
@@
Figure 14: Effect of the proposed attack strategies against the CRH decoder. The simulation parameters are $k$=20 and $S$=60. The legends for the top two subplots follow for the bottom two subplots.
@@
[…]
@@ 4.3 Cost Constrained Optimal Attacks
In the previous section, we have characterized the decoding failure rates of various LT decoders for arbitrary values of $\sigma_{0}$ and $\sigma$ such that $\sigma_{0}\geq\sigma$. However, in practice, there are costs associated with reading and erasing droplet nodes. In this regard, we denote $c_{r}$ and $c_{e}$ as the cost to read and erase one droplet node, respectively. Therefore, we introduce attack-cost as a new metric of interest, which is defined as follows.
@@
In practice, $c_{r}$ is quantified as the time taken by an adversary to contact one droplet node and read its contents and $c_{e}$ is quantified as the time taken to erase one droplet node. Therefore, the attack-cost is quantified as the total time taken by an adversary to launch the attack. Typically, the attack-cost is upper-bounded in practice, to which we propose a method to choose the optimal pair ($\sigma_{0}$,$\sigma$) using which the cost constrained adversary can incur maximum failure rates on the LT decoders.
@@
In practice, given $S$, $\nu$ and $\zeta$, the adversary would need to choose the pair ($\sigma_{0}$,$\sigma$) such that the failure rates of the targeted decoder is maximized under the constraint in (7) along with the constraints $0\leq\sigma_{0}\leq 1$ and $0\leq\sigma\leq\sigma_{0}$. Towards reducing the search space for the above problem without compromising on the failure rates, the following theorem presents tighter bounds on $\sigma$ and $\sigma_{0}$.
@@
[…]
tight_bounds The feasible range for the search space on $\sigma$ and $\sigma_{0}$ is to use $\sigma\in(0,\frac{\nu}{S(1+\zeta)}]$, and then to fix $\sigma_{0}=\frac{\nu-\sigma S\zeta}{S}$ for a given value of $\sigma$.
@@
[…]
@@ 5 Conclusion and Future Scope
In the second part, we have investigated the vulnerabilities of LT-coded blockchains against a wide range of non-oblivious DoS attacks. We have identified that the degree of the droplet nodes, which are easily accessible to the adversary, open doors to the proposed class of attacks. Therefore, while adopting coded blockchains, it is critical to be aware of these threats and take precautions to mitigate them. Furthermore, as an interesting direction for future research, new coded architectures can be proposed for blockchains that enhances its resilience against optimized threats while also providing the storage savings and scalability features.
@@ 6 Acknowledgement
This work was supported by the research grant “Secure Networks and Edge-Computing Hardware for Industry 4.0”, funded by the Ministry of Electronics and Information Technology, New Delhi, India.

### v3

@@ 4.2.2 Attack Strategy against OFG Decoder
A reasonable attack strategy against the OFG decoder involves the following steps:
@@
One possible method to carry out the aforementioned attack strategy is to undertake an exhaustive search to find the combination of droplet nodes that produces a matrix of minimum rank. However, this is a computationally expensive approach as it involves the evaluation of ranks of $\sigma_{0}S\choose\sigma S$ matrices, each with dimensions $(\sigma_{0}-\sigma)S\times k$. To circumvent this problem, we present a computationally feasible algorithm, presented in Algorithm 1, as an alternative approach to exhaustive search. We emphasize that the algorithm may not identify the exact $(\sigma_{0}-\sigma)S$ droplet nodes of minimum rank, however, it does provide a solution with low rank.
@@
10:     Swap($\mathbf{R}[j][\cdot]$$\mathbf{R}[\sigma_{0}S-i+2][\cdot]$);
@@ 4.2.3 Attack Strategy against BRH Decoder
A reasonable attack strategy against the BRH decoder is same as that against the OFG decoder.
@@ 4.2.4 Attack Strategy against CRH Decoder
With the CRH decoder, it is clear that its requirements for successfully decoding of the blockchain are in terms of both singleton droplet nodes as well as rank offered by the residual honest droplet nodes. As a result, putting forward a reasonable attack strategy for this case is non-trivial. However, given that the BP part of the CRH decoder is executed before the OFG part, we believe that depleting singletons in the network would help stalling the decoding process. Thus, with the following strategy, we show that attacks in ?THM? LABEL:BP_Optimal_Attack and ?THM? LABEL:OFG_Optimal_Attack are applicable depending on the value of $\eta_{c}$. Recall that $\eta_{c}$ is the number of blocks of an epoch that must be retrieved by the BP part of the CRH decoder.
@@
For the CRH decoder with $\eta_{c}>0$ , we propose to follow the same attack strategy as that against the BP decoder. However, for the CRH decoder with $\eta_{c}=0$ , we propose to follow the same attack strategy as that against the OFG decoder.
@@ 4.3 Experimental Results for Attack Strategies
In this section, first, we provide details on our experimental setup, and then present the impact of the proposed attack strategies.
4.3.1 Experimental Setup
All our experiments are conducted via simulations on a standalone machine by modelling the following three steps. Step 1: Generation of droplets as per the LT encoding process. Step 2: Execution of the proposed threat models on a subset of LT coded droplets based on the read and write capabilties of the attacker, and Step 3: Reconstruction of the blockchain by the bucket nodes in the network using the coded fragments in the live droplet nodes. In order to generate the experimental results related to decoding failure rates, we use Monte-Carlo simulations [53] by implementing Step 1, Step 2 and Step 3 on MATLAB over an ensemble of $10^{4}$ iterations. In each iteration, to implement Step 1, we fix the LT code parameters for the blockchain, and then generate a certain number of droplets in a random manner using the robust Soliton distribution [17]. For Step 2, we fix the read and write capabilities of the attacker, and randomly choose a certain number of droplets with uniform distribution. Subsequently, depending on the attack strategy chosen by the attacker, we make a subset of the read droplet nodes unavailable for other bucket nodes. We realize this task in our simulation model by erasing the contents of the droplet nodes. Finally, for Step 3, we use the rest of the live droplet nodes to reconstruct the blockchain. Given that the droplet generation process is stochastic, the choice of the droplets read by the attacker is stochastic in each interaction, we declare a decoding failure if a bucket node is unable to reconstruct the entire $k$ blocks of blockchain in each iteration. Overall, by aggregating the events of decoding failures over $10^{4}$ iterations, we empirically compute the decoding failure rate, and plot them as a function of various system parameters in the experimental results. This experimental setup is used to generate all the results in the rest of the paper.
4.3.2 Results
@@
Figure 10: Effectiveness of the proposed degree based attack strategy against a bucket node with BP decoder. The parameters for simulations are kk=20, SS=60, $\sigma\in\{0.1,0.2,0.3,0.4\}$ and $\xi\in\{1.3,1.6,1.9,2.2,2.5\}$.
First, we conduct simulations to characterize the decoding failure rate of the BP decoder for a given $\sigma$, as a function of $\xi$. Our experiments are performed for $k=20$, $S=3k$ such that $\sigma\in\{0.1,0.2,0.3,0.4\}$. For each value of $\sigma$, decoding failure rates are presented with $\xi\in\{1.3,1.6,1.9,2.2,2.5\}$ for two cases: one wherein the droplet nodes that are erased are randomly chosen, and the other wherein droplet nodes are chosen after executing the attack strategy presented in ?THM? LABEL:BP_Optimal_Attack. For each ($\sigma$$\xi$) pair, the obtained failure rates are presented in Fig. 10. Based on the blue bars in Fig. 10, it is clear that for a given $\sigma$, decoding failure rate increases with increase in $\xi$ when the droplet nodes are erased by executing the degree based attack strategy. This behaviour is intuitive since the adversary should get advantage as he reads more droplet nodes. Furthermore, when the erased droplet nodes are randomly chosen, referred to as blind attacks, we observe from the green bars that the failure rate remains constant at different values of $\xi$ for a given $\sigma$.
Figure 11: Comparison of degree based and score based attack strategies against a bucket node with BP decoder. The parameters for simulations are kk=20, SS=60, $\sigma\in\{0.05,0.1,0.15,0.2\}$ and $\sigma_{0}\in\{0.6,0.7,0.8,0.9\}$.
Figure 12: Effectiveness of the proposed attack strategy against a bucket node with OFG decoder. The parameters for the simulations are kk=20, SS=60, $\sigma\in\{0.4,0.45,0.5,0.55\}$ and $\xi\in\{1.2,1.4,1.6,1.8\}$.
Next, we present simulation results demonstrating that the score-based attack strategy for BP decoder, proposed in ?THM? LABEL:BP_Score_Attack is more effective than the one described in ?THM? LABEL:BP_Optimal_Attack, in Fig. 11. The parameters used for simulation are specified in the caption of Fig. 11. By comparing the blue and cyan bars in Fig. 11, it is clear that there is an incremental improvement in the failure rate of score-based attack strategy compared to the degree based strategy. We can also observe that this increment is more significant for lower values of $\sigma$. This is intuitive because, as $\sigma$ approaches $\sigma_{0}$, the set of droplet nodes erased by an adversary starts becoming similar for both the strategies. As a result, their respective failure rates also turn out to be almost identical. However, at small $\sigma$, the score-based attack strategy erases the droplet nodes that are of highest priority to the BP decoder, whereas the degree based strategy may not erase the exact same droplet nodes. This creates a significant difference in the failure rates between the two strategies.
@@
Figure 14: Effect of the proposed attack strategies against the CRH decoder. The simulation parameters are kk=20 and SS=60. The legends for the top two subplots follow for the bottom two subplots.
@@
Finally, Fig. 14 demonstrates the comparison on the effectiveness of the attack strategies given in ?THM? LABEL:BP_Score_Attack and ?THM? LABEL:OFG_Optimal_Attack on the CRH decoder. The experiments are performed for $k=20$, $S=3k$ and $\sigma=0.3$. The failure rates of the CRH decoder corresponding to random erasing (green bars), minimizing rank strategy (red bars) and executing score-based attack strategy (cyan bars) are obtained for values of $\xi\in\{1.5,2,2.5,3\}$. The experiments were conducted for values of $\eta_{c}\in\{1,16\}$, in order to study the behaviour of the employed attack strategies, for near-extreme values of $\eta_{c}$. From the plots, we infer that for any non-zero values of $\eta_{c}$, the attack strategy in ?THM? LABEL:BP_Score_Attack produces higher decoding failure rates. However, for $\eta_{c}=0$, the attack strategy in ?THM? LABEL:OFG_Optimal_Attack maximises the failure rate because the CRH decoder does not necessarily require any singleton for decoding and hence minimizing the rank of honest droplet nodes is sufficient.
@@ 4.4 Cost Constrained Optimal Attacks
In the previous section, we have characterized the decoding failure rates of various LT decoders for arbitrary values of $\sigma_{0}$ and $\sigma$ such that $\sigma_{0}\geq\sigma$. However, in practice, there are costs associated with reading and erasing droplet nodes. In this regard, we denote $c_{r}>0$ and $c_{e}>0$ as costs to read and erase one droplet node, respectively. Therefore, we introduce attack-cost as a new metric of interest, which is defined as follows.
@@
In practice, $c_{r}$ is quantified as the time taken by an adversary to contact one droplet node and read its contents and $c_{e}$ is quantified as the time taken to erase one droplet node. Therefore, the attack-cost is quantified as the total time taken by an adversary to launch the attack. Typically, the attack-cost is upper-bounded in practice, to which we propose a method to choose the optimal pair ($\sigma_{0}$$\sigma$) using which the cost constrained adversary can incur maximum failure rates on the LT decoders.
@@
In practice, given $S$, $\nu$ and $\zeta$, the adversary would need to choose the pair ($\sigma_{0}$$\sigma$) such that the failure rates of the targeted decoder is maximized under the constraint in (15) along with the constraints $0\leq\sigma_{0}\leq 1$ and $0\leq\sigma\leq\sigma_{0}$. Towards reducing the search space for the above problem without compromising on the failure rates, the following theorem presents tighter bounds on $\sigma$ and $\sigma_{0}$.
@@
The feasible range for the search space on $\sigma$ and $\sigma_{0}$ is to use $\sigma\in(0,\frac{\nu}{S(1+\zeta)}]$, and then to fix $\sigma_{0}=\frac{\nu-\sigma S\zeta}{S}$ for a given value of $\sigma$.
@@
Using ?THM? LABEL:tight_bounds, we are able to reduce the search space to one-dimension, thereby reducing the attack complexity. Fig. 15 to Fig. 18 demonstrates the evaluation of decoding failure rates of BP, OFG, BRH and CRH decoders respectively, for values of $\sigma$ in the range $0<\sigma\leq\sigma_{max}(\zeta)$ with a step size of 0.01, and their corresponding maximum values of $\xi$. All four decoders use $k$ and $S$ values of 20 and 60, respectively. Fig. 15 finds the best ($\sigma_{0}$$\sigma$) using the score based attack strategy for BP decoder, wherein the normalised attack-cost of the adversary is constrained to $\nu=1\times S$. Fig. 16 is for the optimal OFG attack, where $\nu=1.5\times S$. Fig. 17 corresponds to failure rates resulting from attacking the BRH decoder with $K=44$ and $\nu=1.4\times S$. Fig. 18 shows failure rates resulting from CRH decoder with $\eta_{c}=6$ and $\nu=1\times S$. Also, we have used $\zeta\in\{1.25,1.5,1.75,2\}$ for all the experiments pertaining to cost constrained attacks. Note that we have used the values of $c_{e}$ and $c_{r}$ such that $c_{e}\geq c_{r}$, i.e., $\zeta\geq 1$. This condition stems from our assumption that the costs incurred for executing a DoS threat on a droplet node is typically much higher than reading the contents of a droplet node. Otherwise, using the values of $\zeta<1$ would imply that the coded blockchain is weak enough to encourage adversarial entities to erase the contents of a droplet node with efforts much lower than that of downloading the contents of a droplet node. This is not a desired feature of a secure coded blockchain network.
@@ 5 Conclusion and Future Scope
In the second part, we have investigated the vulnerabilities of LT-coded blockchains against a wide range of non-oblivious DoS attacks. We have identified that the degree of the droplet nodes, which are easily accessible to the adversary, open doors to the proposed class of attacks. The attacks proposed in Section 4 have carefully leveraged on the underlying algebraic structure of the BP and the OFG decoders of LT codes. While we were successful in convincing that these strategies make a significant impact on scalability, further studies on optimal attack strategies under the class of non-oblivious attacks are interesting directions for future research. When adopting coded blockchains, it is critical to be aware of these threats and take precautions to mitigate them. Furthermore, as an interesting direction for future research, new coded architectures can be proposed for blockchains that enhances its resilience against optimized threats while also providing the storage savings and scalability features. The novelty of our threat model lies in the choice of the droplet nodes for executing the DoS attack by the attacker. Therefore, we used simulation based experiments to validate the impact of the proposed contributions. However, if one is interested in measuring other peripheral metrics such as delays associated with implementing the attacks, and the impact of synchronization issues between the stages of attack and blockchain reconstruction, then these attacks could also be implemented on a hardware setup over a network of servers. These are interesting directions for future research.
@@ 6 Acknowledgement
This work was supported by the research grant “Secure Networks and Edge-Computing Hardware for Industry 4.0,” funded by the Ministry of Electronics and Information Technology, New Delhi, India.